# Confidence-Weighted Boundary-Aware Learning for Semi-Supervised Semantic Segmentation

### v1

@@ IV-C Quantitative Analysis
Figure 4: Qualitative comparisons of CW-BASS with other state-of-the-art methods, ST++ [9] and UniMatch [36] on the PASCAL VOC 2012 and Cityscapes datasets. All methods are compared using ResNet-50 Backbone under the 1/8 setting for validation.
IV-C1 Comparison on Computational Costs and Training Flows
Figure 5 illustrates the convergence behavior and performance of CW-BASS compared to the self-training method ST++ [9] on the Cityscapes dataset during the initial 20 epochs, utilizing the 1/16 partition protocol. CW-BASS achieves faster stabilization due to its confidence-weighted loss, which prioritizes high-confidence predictions and mitigates the influence of noisy pseudo-labels. Furthermore, its boundary-aware learning enhances segmentation accuracy near object edges, contributing to more rapid convergence. In contrast, ST++ depends on iterative retraining after selecting reliable images post pseudo-labeling. Overall, CW-BASS integrates swift initial convergence with sustained performance improvements throughout the training process.
Table II summarises data utilisation and training overhead for various approaches, highlighting differences in the number of networks trained, unlabeled data processed per epoch, and total training epochs. CPS has a dual network design and treats the labeled set, $\mathcal{D}_{L}$ without the ground truth as additional unlabeled set, increasing the volume of training data. On the other hand, PS-MT implements consistency learning with perturbations to both teacher and student networks. CPS trains two networks, processing 10.5k and 2.9k unlabeled images per epoch on Pascal VOC and Cityscapes, respectively. PS-MT employs three networks with aggressive augmentations, often doubling or tripling unlabeled data (e.g., 19.8k samples per epoch for the 1/16 Pascal partition) and requires 320–550 epochs on Cityscapes. ST++ trains four networks with slightly fewer unlabeled samples per epoch compared to PS-MT (14.8k on Pascal and 4.1k on Cityscapes, both below 1/16). In contrast, CW-BASS trains two networks with fewer augmentations, using the least unlabeled data (up to 9.9k on Pascal and 2.7k on Cityscapes) and converging in just 80 epochs on Pascal and 240 epochs on Cityscapes, regardless of partition. This balanced approach achieves competitive performance while avoiding the overheads of intensive augmentations or large network ensembles.
IV-C2 Comparison on Segmentation Performance
Results on Pascal VOC. Table I compares performance on the Pascal VOC 2012 dataset using various state-of-the-art (SOTA) semi-supervised semantic segmentation methods. Unlike approaches that use multiple teacher networks and various perturbations, our method relies on a single model for self-training. To highlight our performance, we also specifically compare our results with ST++, an advanced and simple self-training approach.
Using only 1/8 (12.5%) of the labeled training data, our method produces a mean Intersection over Union (mIoU) of 75.81%, outperforming existing state-of-the-art approaches. This also exceeds the performance of the fully supervised baseline by a remarkable 11.2%, demonstrating the effectiveness of our semi-supervised learning method in limited label settings.
IV-D Qualitative Analysis
Fig. 1 and Fig. 4 show visual segmentation outputs produced by our method, CW-BASS on Pascal VOC and Cityscapes dataset, compared against a prominent self-training state-of-the-art method, ST++ [9] and UniMatch [36]. The qualitative gains are shown, with our method consistently providing clearer object boundaries, more accurate region delineations, and better overall segmentation performance. These results reflect the quantitative findings, showing the robustness and versatility of the proposed framework.
@@ IV-E1 Ablation Analysis
Confidence-Weighted Loss ($L_{labeled}$ + $L_{weighted}$) Confidence-Weighting serves as the main component in our method ($L_{\text{labeled}}$) and ($L_{\text{weighted}}$) and achieves a mean Intersection over Union (mIoU) of 73.43%. This result is much higher than the baseline and some previous SOTA methods, underscoring the importance of the weighted loss in prioritizing high-confidence pseudo-labels and effectively mitigate the impact of label noise on model performance.
Boundary-Aware Loss ($L_{boundary}$). Incorporating boundary-aware loss into the confidence-weighting, significantly improves segmentation performance, particularly near object boundaries. This addition results in an mIoU of 74.67%, marking a notable improvement of 6.36%. The findings demonstrate the effectiveness of addressing boundary blur through specialized loss functions.
Dynamic Thresholding Mechanism ($T$). The integration of a dynamic thresholding mechanism enhances the model’s ability to filter out low-confidence pseudo-labels. Using this mechanism in conjunction with $L_{\text{labeled}}$ and $L_{\text{weighted}}$ yields an mIoU of 69.01%, highlighting its role in dynamically refining pseudo-label quality by discarding unreliable predictions.
Confidence Decay Strategy ($\alpha$). The confidence decay strategy gradually reduces the influence of low-confidence pseudo-labels during training. This approach stabilizes learning by improving pseudo-label reliability over time. The configuration ($L_{\text{labeled}}+L_{\text{weighted}}+\alpha$) produces modest improvements, showcasing its ability to complement other strategies by enhancing pseudo-label robustness.
Discussion. The ablation study highlights the Confidence-Weighted Loss and Boundary-Aware Loss as the primary contributors to performance gains. The significant improvement observed with boundary-aware loss highlights the effectiveness of integrating boundary-delineation techniques in semi- supervised frameworks. The Confidence Decay Strategy and Dynamic Thresholding Mechanism address the challenges of noisy pseudo-labels and confirmation bias.
@@ IV-E2 Hyperparameter Analysis
To investigate the impact of our hyperparameter on model performance, we perform experiments using two groups of settings: our standard & default setting (i.e., $\gamma=1.0,\beta=0.5,\alpha=0.85$), which we use for all our experiments, and a more conservative setting ($\gamma=0.5,\beta=1.0,\alpha=1.0$), as shown in Table IV. The standard settings batch yield the best model performance because the parameters are balanced, whereas the conservative settings yield an mIoU 4.41% lower than the standard settings.
For Confidence weighting, a $\gamma$ value of 0.5 reduces the influence of these labels, making the model more conservative to minimize the risk of overfitting to incorrect labels. On the other hand, setting ($\gamma$) to 2.0 increases the emphasis on high-confidence pseudo-labels, encouraging the model to reinforce confident predictions more aggressively.
@@
Finally, the confidence decay factor ($\alpha$), with a default value of 0.9, can be adjusted for specific training needs. Lowering $\beta$ to 0.85 accelerates the decay, which is particularly useful when a more aggressive reduction of noisy pseudo-labels is required.
@@ V Conclusion
In this work, we presented Confidence-Weighted Boundary-Aware Semantic Segmentation (CW-BASS), a novel framework for semi-supervised semantic segmentation that addressed key challenges such as confirmation bias, boundary blur, and label noise. CW-BASS integrates confidence-weighting, dynamic thresholding, and boundary-delineation techniques to achieve state-of-the-art performance. CW-BASS significantly reduces the computational cost by using only two networks, fewer augmentations, and fewer unlabeled data samples per epoch, resulting in fewer training epochs while maintaining competitive performance. This efficient design is especially useful for resource-constrained scenarios because it balances performance and cost-effectiveness. Furthermore, CW-BASS’s fast initial convergence and robust boundary delineation make it suitable for real-world applications with limited computational resources and unlabeled data.

### v2

@@ IV-E1 Ablation Analysis
Baseline ($L_{labeled}$). We follow standard SSSS practice by using $L_{labeled}$, also referred to as SupOnly (Supervised Only), as the baseline for our framework.
Confidence-Weighted Loss ($L_{labeled}$ + $L_{weighted}$). Confidence-weighting is central to our method, combining $L_{\text{labeled}}$ and $L_{\text{weighted}}$, and achieves an mIoU of 73.43%. This significantly outperforms the baseline and some SOTA methods, showing how prioritizing high-confidence pseudo-labels mitigates label noise and boosts performance.
Boundary-Aware Loss ($L_{boundary}$). Adding boundary-aware loss to confidence-weighting improves segmentation near object edges, yielding an mIoU of 74.67%;a 6.36% gain over the baseline. This confirms the effectiveness of handling boundary blur with specialized loss functions.
Dynamic Thresholding Mechanism ($T$). Dynamic thresholding helps filter out low-confidence pseudo-labels. When combined with $L_{\text{labeled}}$ and $L_{\text{weighted}}$, it achieves an mIoU of 69.01%, demonstrating its role in refining pseudo-label quality during training.
Confidence Decay Strategy ($\alpha$). This strategy gradually reduces the influence of uncertain pseudo-labels, stabilizing learning over time. The setup ($L_{\text{labeled}}+L_{\text{weighted}}+\alpha$) offers modest gains, showing its value in enhancing pseudo-label robustness.
Discussion. The study identifies Confidence-Weighted Loss and Boundary-Aware Loss as key contributors to performance. The substantial gain from boundary-aware loss highlights the value of integrating boundary-delineation. Meanwhile, Dynamic Thresholding and Confidence Decay effectively address noisy pseudo-labels and confirmation bias.
@@ IV-E2 Hyperparameter Analysis
To investigate the impact of our hyperparameter on model performance, we perform experiments using two groups of settings: our standard & default setting (i.e., $\gamma=1.0,\beta=0.5,\alpha=0.85$), which we use for all our experiments, and a more conservative setting ($\gamma=0.5,\beta=1.0,\alpha=1.0$), as shown in Table IV. The standard settings batch yield the best model performance because the parameters are balanced, whereas the conservative settings yield a lower mIoU than the standard settings.
For Confidence weighting, $\gamma$ = 0.5, 1.0, 2.0 was tested. A $\gamma$ value of 0.5 reduces the influence of these labels, making the model more conservative to minimize the risk of overfitting to incorrect labels. A $\gamma$ value of 1.0 achieves optimal performance whilst setting $\gamma$ to 2.0 increases the emphasis on high-confidence pseudo-labels resulting in the model to reinforcing confident predictions more aggressively.
@@
Finally, the confidence decay factor ($\alpha$), with a default value of 0.9, can be adjusted for specific training needs. Lowering $\beta$ to 0.85 accelerates the decay, which could be particularly useful when a more aggressive reduction of noisy pseudo-labels is required.
@@ V Conclusion
In this work, we introduced Confidence-Weighted Boundary-Aware Semantic Segmentation (CW-BASS), a new SSSS framework that tackles key challenges including confirmation bias, boundary blur, and label noise to achieve state-of-the-art performance. It reduces computational cost by using only two networks, fewer augmentations, and less unlabeled data per epoch, leading to faster convergence with competitive results. This efficient design is well-suited for resource-constrained settings, balancing performance and cost. Additionally, CW-BASS’s convergence and strong boundary handling make it practical for real-world applications with limited data and compute.
For future work, our method could explore transformer-based backbones like DINO-ViT to enhance representation learning and pseudo-label reliability in sparse-label scenarios. Broader evaluations on datasets such as COCO and ADE20K would further validate its generalization across diverse, complex scenes.